For a solid A, the changes in volume and entropy at melting are given as 3.8x10-6 m3/mol and 24 j mol-1 K-1. How much pressure should be applied to increase the melting point by 2.5°?
Answer #1

Answer : we know that i Following are equations, du= Tds - Pdv (I) = - ( 2 ) dH = Tas +vap = (ST) = ( 2 ) & we are given, avChange în pressure = ap op= 24 x 25 Imoll kalk mm mol 3.8 x 10-6 Im3 dp = 15*789*10° unit of Energy a joule, dimension = M 2